$SPAD/src/input d02kef.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
xpoint:Matrix SF:=
 [[0.0 ,0.1 ,4**(1/3) ,30.0 ,30.0 ]]
m:=5
k:=11
tol:SF:=0.0001
maxfun:=0
match:=0
elam:=14
delam:=1
hmax:Matrix SF:=
 [[0.0 ,0.0 ],[0.0 ,0.0 ],_
  [ 0.0 ,0.0 ],[0.0 ,0.0 ],[0.0 ,0.0 ]]
maxit:= 0
vef:Vector Expression Float:=
 [1.0 ,ELAM-X-2.0/(X*X) ,1.0 ]
coeffn:Asp10(COEFFN):=retract vef
mef:Matrix Expression Float:=
 [[XL ,2.0 ],[1.0 ,-sqrt(XR-ELAM) ]]
bdyval:Asp80(BDYVAL):=retract mef
result:=d02kef(xpoint,m,k,tol,maxfun,match,elam,delam,hmax,maxit,-1,coeffn,bdyval)
-- the following are the default Asps used.
outputAsFortran()$Asp12(MONIT)
outputAsFortran()$Asp33(REPORT)

———-

References^{[1] nothing}

4